# Equivalence of scalar-tensor theories and scale-dependent gravity

### v1

@@ III.3 Equivalence on the level of field equations
It is known that the equivalences between $f(R)$-gravity and SD gravity or STTs on the level of the actions automatically imply equivalences on the level of field equations Calzada et al. (2023); Olmo (2007). Therefore, we will now investigate the equivalence between SD gravity and STTs on the level of the field equations as another consistency check. This also gives us the opportunity to further analyze the two classes of theories, and this will uncover an important relation between the EOM for $\phi$ and a specific scale-setting method used in SD gravity.
For SD gravity and STTs, we will need to derive the Jordan frame Einstein equations, which are modified not only due to the presence of the kinetic terms for $k$ and $\phi$, but also due to the non-minimal coupling of the Ricci scalar to $\frac{1}{G(k)}$ and $\frac{1}{A^{2}(\phi)}$. In addition, for STTs, we will have an EOM for $\phi$. Note that, for SD gravity, the action in Eq. (7) is not varied with respect to $G$ and $\Lambda$ because the running couplings are not obtained from a variational principle but rather from solving the renormalization group equations once a scale-setting relation $k(x)$ has been fixed Reuter and Weyer (2004a). The couplings can then be seen as externally prescribed background fields in the context of the variational principle. Since, as of yet, there is no unique way of fixing $k$ and various methods are discussed in the literature Bertini (2025); Koch et al. (2015); Contreras et al. (2016); Domazet and Stefancic (2012), the focus is often not on finding $k(x)$ itself, but rather on determining the functions $G(k(x))$ and $\Lambda(k(x))$. One method is the so-called effective Lagrangian fixed point condition (ELFP) Koch et al. (2015); Contreras et al. (2016), which reads
@@
which coincides with the EOM for $k$ (49) multiplied by $2\nabla_{\nu}k$. Thus, imposing Eq. (49) is a sufficient condition for the vanishing of the divergence of the effective energy-momentum tensor. Note that working with Eq. (III.3) instead of Eq. (50) is merely a matter of preference since we can arrive at the same conclusion by working with either of them. Performing a similar calculation for STTs yields the same result, i.e., the appropriate analogue of Eq. (III.3) is proportional to the EOM for $\phi$ (III.3) and thus vanishes on-shell. So, while STTs are always anomaly-free on-shell, for SD gravity we need the additional constraint (49). In particular, note that Eq. (49) is valid only for the specific action considered here. For effective actions that include additional ingredients, such as an electromagnetic source, for instance, the condition Eq. (49) must be generalized to incorporate the running of the electromagnetic coupling function. Ref. Reuter and Weyer (2004a) analyzed the divergence of the effective energy-momentum tensor of SD gravity with an additional contribution to the energy-momentum tensor describing the four-momentum carried by the running couplings. Since Ref. Reuter and Weyer (2004a) did not treat $k$ as a dynamical field, no explicit kinetic term for $k$ was included in the action and no EOM for $k$ was imposed. Instead, in order to guarantee the vanishing divergence of the energy-momentum tensor, suitable choices of the additional contribution to the energy-momentum tensor, the matter-part of the energy-momentum tensor, as well as the cosmological constant, were discussed. One particular example was $G(k)\Lambda(k)\!\!:\text{const.}$ and $T_{\mu\nu}=0$, where $T_{\mu\nu}$ is the matter part of the energy-momentum tensor. In this case, it was found that the additional contribution has to be
@@
we obtain $\theta_{\mu\nu}=\Delta_{B}t_{\mu\nu}$. Moreover, Ref. Reuter and Weyer (2004a) found that $\theta_{\mu\nu}$ can arise from a term in the action
@@
which, up to a numerical prefactor that we can leave out, coincides with the kinetic term in Eq. (7) for the choice of $B(k)$ made in Eq. (54). Thus, we have found a kinetic term for $k$, such that the situation in Ref. Reuter and Weyer (2004a) can be treated as a special case of our discussion for the divergence of the energy-momentum tensor corresponding to the action in Eq. (7), in which the particular choices (54) and $G(k)\Lambda(k)\!\!:\text{const.}$ lead to a vanishing divergence of the energy-momentum tensor without the need to impose the EOM for $k$.
@@ III.4 Equivalence in the Einstein frame
In Sec. III.3, we have seen that the EOMs of SD gravity and STTs in the Jordan frame are rather complicated and coupled to the Einstein equations, which renders the search for solutions intricate. However, when dealing with concrete models, we want to give explicit solutions for the EOMs. From STTs we know that the EOMs in the Einstein frame are much simpler than in the Jordan frame; see, e.g., Ref. Pitschmann (2023). Furthermore, there are no couplings to the Einstein equations in this frame. Therefore, we will now briefly discuss the Einstein frame version of the equivalence between SD gravity and STTs.
@@
Given Eq. (69), the EOM (76) appears to be rather complicated. However, it can be simplified when considering a SD theory with $B(k)=0$, which applies to essentially all models in the literature Rincón (2019); Alvarez et al. (2024a, 2022, b); Rincon et al. (2023). Furthermore, in most of these models, it is not required that $k$ satisfies its EOM. In such cases, we only need to impose the EOM for $\phi$. However, as we have seen in Sec. III.3, it might not always be possible for $\phi$ to satisfy its EOM if the one for $k$ is not imposed. Another option would be to start from a STT and make the choice $k(\phi)=\phi$; see Sec. III.1.2. In this case, Eq. (70) reduces to $\bar{B}(k)=1/2$, such that Eq. (76) becomes
@@ IV.1 Consistency conditions for physical models
Other important aspects are the signs of the kinetic terms for $\phi$ and $k$. It is known that a kinetic term with the wrong sign in the action leads to ghost-like instabilities of the corresponding theory Delhom et al. (2022). From Eq. (10), we can immediately see that for both directions of the equivalence the kinetic term of the resulting theory will have the same sign as the original one. Thus, starting with a theory that is free of instabilities, this property will be inherited by the equivalent theory.
@@ IV.2 STTs as SD theories
Here, we will apply the equivalence to three popular STTs: conformally coupling chameleons, symmetrons and environment-dependent dilatons. For this, we will follow the presentation in Ref. Fischer et al. (2024c), and we will work in the Einstein frame. In the Einstein frame, the conformal coupling to the metric tensor induces an effective coupling to matter, see, for example, Ref. Käding (2019), such that the scalar $\phi$ has an effective potential
@@
where $\rho(x)$ is the density of non-relativistic matter. The matter coupling is expected to cause a fifth force, which, however, is screened in dense environments due to the interplay with $V(\phi)$, such that tight Solar System-based constraints on fifth forces Dickey et al. (1994); Adelberger et al. (2003); Kapner et al. (2007) can be circumvented. Since, in Sec. III.4, we have shown that SD gravity can also be expressed in the Einstein frame, we would expect that the conformal coupling of the form given in Eq. (66) will also lead to an additional coupling to matter, whose significance could be explored in future works. The scalar potential and scale factor that define each STT are listed for the three considered models in Tab. 1.
@@
Table 1: Taken from Ref. Fischer et al. (2024c); list of the quantities defining the three considered STTs; Chameleon: different models are distinguished by $n\in\mathbb{Z}^{+}\cup 2\mathbb{Z}^{-}\setminus\{-2\}$. $\Lambda$ and $M_{C}$ are mass scales that describe the self-interaction of the chameleon and the coupling to matter, respectively. Symmetron: The tachyonic mass $\mu$ and the dimensionless self-coupling parameter $\lambda_{S}$ determine the symmetron potential, while the mass scale $M_{S}$ sets the matter coupling. Dilaton: A constant energy density $V_{0}$ and a dimensionless self-coupling constant $\lambda$ describe the dilaton potential, and the dimensionless parameter $A_{2}$ determines the coupling to matter.
@@
The form of $k(x)$ is determined by the specific solution of the EOM for $\phi$. For example, the solution for a chameleon around a static, homogeneous sphere with radius $\mathcal{R}$ is given by Khoury and Weltman (2004a)
@@
For symmetrons and environment-dependent dilatons, the solutions for $\phi$ around a static, homogeneous sphere are qualitatively similar to Eq. (81); see Refs. Burrage et al. (2016); Brax et al. (2022). Consequently, for this particular setup, and from Tab. 2, we conclude that, in both cases, $G(k)$, $\Lambda(k)$ and $k(x)$ are smooth, $G$ is always larger than 0, and $k$ falls off exponentially fast, reaching a constant value at infinity. Furthermore, for both models, the scalar potentials and the resulting $\Lambda(k)$ are bounded from below. Therefore, we have shown that also symmetrons and dilatons lead to consistent equivalent SD gravities. Note that the results will likely be very different if another choice for the relation between $\phi(x)$ and $k(x)$ is made.
@@ IV.3 SD gravity models as STTs
Now, we will also demonstrate how the equivalence relations are applied to various SD models from the literature. In contrast to Sec. IV.2, where we have worked in the Einstein frame, we will now work in the Jordan frame. Since the models discussed in the literature do not feature a kinetic term for $k$ Rincón (2019); Alvarez et al. (2024a, 2022, b); Rincon et al. (2023), we have $B(k)=0$ in all examples, which means that the resulting STTS will all have scale factors of the form given in Eq. (16) and only differ by their potentials.
@@ IV.3.1 Einstein-Hilbert cosmology with NEC
As a first example, we will discuss SD Einstein-Hilbert cosmology in the absence of matter; see, e.g., Refs. Alvarez et al. (2024a, 2022, 2021); Hipólito-Ricaldi et al. (2025); Bertini et al. (2024). We assume a homogeneous and isotropic spacetime that is spatially flat. In this case, the only relevant coordinate is the cosmological time $t$ and we can directly work with time-dependent gravitational couplings. The modified Friedmann equations following from Eq. (41) are
@@
The solution to the combined system of equations is
@@
to which we can apply Eqs. (12) and (13) to obtain the quantities of the equivalent STT:
@@ IV.3.2 Einstein-Hilbert cosmology with ELFP
As a second example, we investigate the SD bouncing cosmology model of Ref. Alvarez et al. (2024b). Again, the only relevant coordinate is the cosmological time $t$ and we will directly work with time-dependent gravitational couplings. In bouncing cosmologies, the NEC is automatically violated, which is why the ELFP is employed to uniquely fix the theory. It is important to note that the condition imposed in Ref. Alvarez et al. (2024b) is actually stronger than the ELFP given by Eq. (40). More precisely, it is demanded that Eq. (40) is satisfied for any $k(t)$, which is only possible for a constant Lagrangian. In this situation, the EOM for $k$ trivializes. Since the Lagrangian of the equivalent STT will also be constant, the EOM for $\phi$ is trivial as well and no inconsistencies of the equivalence arise at the level of the field equations. Again using Eqs. (83) and (84), we complement them by the stronger version of the ELFP:
@@
and $c_{1}=-t_{b}$, $c_{R}=\tilde{R}_{0}=4\Lambda_{0}$ and $c_{2}=\sqrt{2}a_{0}e^{\sqrt{\frac{\Lambda_{0}}{3}}t_{b}}$ with $t_{b}$ being the bouncing time, $\tilde{R}_{0}$ being the constant Ricci scalar, which is related to the observed value of the cosmological constant $\Lambda_{0}$, and $a_{0}$ being related to the scale factor. $G_{i}$ is another integration constant. For more details we refer to Ref. Alvarez et al. (2024b). Due to the complicated form of $G(t)$, we refrain from writing it out explicitly from now on. For this model, Eqs. (12), (13) and (18) give
@@
We will now apply the discussion in Ref. Alvarez et al. (2024b) of the functional behavior of $G(t)$ to the equivalent STT. While the specific properties of $G$ crucially depend on the parameter choice, there are some general features that can be interpreted in the present context. Most importantly, at the bouncing time $t=t_{b}$, $G(t)$ actually diverges towards positive infinity from the right and to negative infinity from the left. The latter violates our requirement that $G(t)>0$, our assumption about smoothness as well as the condition $B(k)>-\frac{3}{4}\frac{1}{G(k)^{3}}\left(\frac{\partial G(k)}{\partial k}\right)^{2}$. Due to the dependence of $\phi$, $A^{2}$ and $V$ on $G$, this affects the whole STT. The scalar potential cannot be bounded from below, and, since $G(t)$ falls off to zero for very early and late times, $\phi$ diverges at infinity. Even worse, $\phi$ is given as the logarithm of $G$, which is not a real number for negative arguments. Overall, while the ELFP indeed allows us to arrive at a STT in which $\phi$ satisfies its EOM, the criteria of the equivalence are violated and it is not possible to define a consistent STT corresponding to this particular SD model.
@@ IV.3.3 Polytropic black hole with NEC
In the last example, we will deal with the polytropic black hole solution obtained in Ref. Contreras et al. (2018). Due to the fact that this is a static, spherically symmetric solution, the only relevant variable is the radius $r$. Applying the NEC (79) to the effective energy-momentum tensor of SD gravity gives

### v2

@@ III.3 Equivalence on the level of field equations
It is known that the equivalences between $f(R)$-gravity and SD gravity or STTs on the level of the actions automatically imply equivalences on the level of field equations [111, 128]. Therefore, we will now investigate the equivalence between SD gravity and STTs on the level of the field equations as another consistency check. This also gives us the opportunity to further analyze the two classes of theories, and this will uncover an important relation between the EOM for $\phi$ and a specific scale-setting method used in SD gravity.
For SD gravity and STTs, we will need to derive the Jordan frame Einstein equations, which are modified not only due to the presence of the kinetic terms for $k$ and $\phi$, but also due to the non-minimal coupling of the Ricci scalar to $\frac{1}{G(k)}$ and $\frac{1}{A^{2}(\phi)}$. In addition, for STTs, we will have an EOM for $\phi$. Note that, for SD gravity, the action in Eq. (7) is not varied with respect to $G$ and $\Lambda$ because the running couplings are not obtained from a variational principle but rather from solving the renormalization group equations once a scale-setting relation $k(x)$ has been fixed [99]. The couplings can then be seen as externally prescribed background fields in the context of the variational principle. Since, as of yet, there is no unique way of fixing $k$ and various methods are discussed in the literature [129, 104, 105, 130], the focus is often not on finding $k(x)$ itself, but rather on determining the functions $G(k(x))$ and $\Lambda(k(x))$. One method is the so-called effective Lagrangian fixed point condition (ELFP) [104, 105], which reads
@@
which coincides with the EOM for $k$ (49) multiplied by $2\nabla_{\nu}k$. Thus, imposing Eq. (49) is a sufficient condition for the vanishing of the divergence of the effective energy-momentum tensor. Note that working with Eq. (III.3) instead of Eq. (50) is merely a matter of preference since we can arrive at the same conclusion by working with either of them. Performing a similar calculation for STTs yields the same result, i.e., the appropriate analogue of Eq. (III.3) is proportional to the EOM for $\phi$ (III.3) and thus vanishes on-shell. So, while STTs are always anomaly-free on-shell, for SD gravity we need the additional constraint (49). In particular, note that Eq. (49) is valid only for the specific action considered here. For effective actions that include additional ingredients, such as an electromagnetic source, for instance, the condition Eq. (49) must be generalized to incorporate the running of the electromagnetic coupling function. Ref. [99] analyzed the divergence of the effective energy-momentum tensor of SD gravity with an additional contribution to the energy-momentum tensor describing the four-momentum carried by the running couplings. Since Ref. [99] did not treat $k$ as a dynamical field, no explicit kinetic term for $k$ was included in the action and no EOM for $k$ was imposed. Instead, in order to guarantee the vanishing divergence of the energy-momentum tensor, suitable choices of the additional contribution to the energy-momentum tensor, the matter-part of the energy-momentum tensor, as well as the cosmological constant, were discussed. One particular example was $G(k)\Lambda(k)\!\!:\text{const.}$ and $T_{\mu\nu}=0$, where $T_{\mu\nu}$ is the matter part of the energy-momentum tensor. In this case, it was found that the additional contribution has to be
@@
we obtain $\theta_{\mu\nu}=\Delta_{B}t_{\mu\nu}$. Moreover, Ref. [99] found that $\theta_{\mu\nu}$ can arise from a term in the action
@@
which, up to a numerical prefactor that we can leave out, coincides with the kinetic term in Eq. (7) for the choice of $B(k)$ made in Eq. (54). Thus, we have found a kinetic term for $k$, such that the situation in Ref. [99] can be treated as a special case of our discussion for the divergence of the energy-momentum tensor corresponding to the action in Eq. (7), in which the particular choices (54) and $G(k)\Lambda(k)\!\!:\text{const.}$ lead to a vanishing divergence of the energy-momentum tensor without the need to impose the EOM for $k$.
@@ III.4 Equivalence in the Einstein frame
In Sec. III.3, we have seen that the EOMs of SD gravity and STTs in the Jordan frame are rather complicated and coupled to the Einstein equations, which renders the search for solutions intricate. However, when dealing with concrete models, we want to give explicit solutions for the EOMs. From STTs we know that the EOMs in the Einstein frame are much simpler than in the Jordan frame; see, e.g., Ref. [118]. Furthermore, there are no couplings to the Einstein equations in this frame. Therefore, we will now briefly discuss the Einstein frame version of the equivalence between SD gravity and STTs.
@@
Given Eq. (69), the EOM (76) appears to be rather complicated. However, it can be simplified when considering a SD theory with $B(k)=0$, which applies to essentially all models in the literature [123, 124, 125, 126, 119]. Furthermore, in most of these models, it is not required that $k$ satisfies its EOM. In such cases, we only need to impose the EOM for $\phi$. However, as we have seen in Sec. III.3, it might not always be possible for $\phi$ to satisfy its EOM if the one for $k$ is not imposed. Another option would be to start from a STT and make the choice $k(\phi)=\phi$; see Sec. III.1.2. In this case, Eq. (70) reduces to $\bar{B}(k)=1/2$, such that Eq. (76) becomes
@@ IV.1 Consistency conditions for physical models
Other important aspects are the signs of the kinetic terms for $\phi$ and $k$. It is known that a kinetic term with the wrong sign in the action leads to ghost-like instabilities of the corresponding theory [131]. From Eq. (10), we can immediately see that for both directions of the equivalence the kinetic term of the resulting theory will have the same sign as the original one. Thus, starting with a theory that is free of instabilities, this property will be inherited by the equivalent theory.
@@ IV.2 STTs as SD theories
Here, we will apply the equivalence to three popular STTs: conformally coupling chameleons, symmetrons and environment-dependent dilatons. For this, we will follow the presentation in Ref. [93], and we will work in the Einstein frame. In the Einstein frame, the conformal coupling to the metric tensor induces an effective coupling to matter, see, for example, Ref. [127], such that the scalar $\phi$ has an effective potential
@@
where $\rho(x)$ is the density of non-relativistic matter. The matter coupling is expected to cause a fifth force, which, however, is screened in dense environments due to the interplay with $V(\phi)$, such that tight Solar System-based constraints on fifth forces [5, 6, 7] can be circumvented. Since, in Sec. III.4, we have shown that SD gravity can also be expressed in the Einstein frame, we would expect that the conformal coupling of the form given in Eq. (66) will also lead to an additional coupling to matter, whose significance could be explored in future works. The scalar potential and scale factor that define each STT are listed for the three considered models in Tab. 1.
@@
Table 1: Taken from Ref. [93]; list of the quantities defining the three considered STTs; Chameleon: different models are distinguished by $n\in\mathbb{Z}^{+}\cup 2\mathbb{Z}^{-}\setminus\{-2\}$. $\Lambda$ and $M_{C}$ are mass scales that describe the self-interaction of the chameleon and the coupling to matter, respectively. Symmetron: The tachyonic mass $\mu$ and the dimensionless self-coupling parameter $\lambda_{S}$ determine the symmetron potential, while the mass scale $M_{S}$ sets the matter coupling. Dilaton: A constant energy density $V_{0}$ and a dimensionless self-coupling constant $\lambda$ describe the dilaton potential, and the dimensionless parameter $A_{2}$ determines the coupling to matter.
@@
The form of $k(x)$ is determined by the specific solution of the EOM for $\phi$. For example, the solution for a chameleon around a static, homogeneous sphere with radius $\mathcal{R}$ is given by [8]
@@
For symmetrons and environment-dependent dilatons, the solutions for $\phi$ around a static, homogeneous sphere are qualitatively similar to Eq. (81); see Refs. [34, 22]. Consequently, for this particular setup, and from Tab. 2, we conclude that, in both cases, $G(k)$, $\Lambda(k)$ and $k(x)$ are smooth, $G$ is always larger than 0, and $k$ falls off exponentially fast, reaching a constant value at infinity. Furthermore, for both models, the scalar potentials and the resulting $\Lambda(k)$ are bounded from below. Therefore, we have shown that also symmetrons and dilatons lead to consistent equivalent SD gravities. Note that the results will likely be very different if another choice for the relation between $\phi(x)$ and $k(x)$ is made.
@@ IV.3 SD gravity models as STTs
Now, we will also demonstrate how the equivalence relations are applied to various SD models from the literature. In contrast to Sec. IV.2, where we have worked in the Einstein frame, we will now work in the Jordan frame. Since the models discussed in the literature do not feature a kinetic term for $k$ [123, 124, 125, 126, 119], we have $B(k)=0$ in all examples, which means that the resulting STTS will all have scale factors of the form given in Eq. (16) and only differ by their potentials.
@@ IV.3.1 Einstein-Hilbert cosmology with NEC
As a first example, we will discuss SD Einstein-Hilbert cosmology in the absence of matter; see, e.g., Refs. [132, 124, 125, 133, 134, 135]. We assume a homogeneous and isotropic spacetime that is spatially flat. In this case, the only relevant coordinate is the cosmological time $t$ and we can directly work with time-dependent gravitational couplings. The modified Friedmann equations following from Eq. (41) are
@@
The solution to the combined system of equations is [132]
@@
with $t_{0}=\pm\sqrt{\frac{3}{\Lambda_{0}}}$, to which we can apply Eqs. (12) and (13) to obtain the quantities of the equivalent STT:
@@ IV.3.2 Einstein-Hilbert cosmology with ELFP
As a second example, we investigate the SD bouncing cosmology model of Ref. [126]. Again, the only relevant coordinate is the cosmological time $t$ and we will directly work with time-dependent gravitational couplings. In bouncing cosmologies, the NEC is automatically violated, which is why the ELFP is employed to uniquely fix the theory. It is important to note that the condition imposed in Ref. [126] is actually stronger than the ELFP given by Eq. (40). More precisely, it is demanded that Eq. (40) is satisfied for any $k(t)$, which is only possible for a constant Lagrangian. In this situation, the EOM for $k$ trivializes. Since the Lagrangian of the equivalent STT will also be constant, the EOM for $\phi$ is trivial as well and no inconsistencies of the equivalence arise at the level of the field equations. Again using Eqs. (83) and (84), we complement them by the stronger version of the ELFP:
@@
and $c_{1}=-t_{b}$, $c_{R}=\tilde{R}_{0}=4\Lambda_{0}$ and $c_{2}=\sqrt{2}a_{0}e^{\sqrt{\frac{\Lambda_{0}}{3}}t_{b}}$ with $t_{b}$ being the bouncing time, $\tilde{R}_{0}$ being the constant Ricci scalar, which is related to the observed value of the cosmological constant $\Lambda_{0}$, and $a_{0}$ being related to the scale factor. $G_{i}$ is another integration constant. For more details we refer to Ref. [126]. Due to the complicated form of $G(t)$, we refrain from writing it out explicitly from now on. For this model, Eqs. (12), (13) and (18) give
@@
We will now apply the discussion in Ref. [126] of the functional behavior of $G(t)$ to the equivalent STT. While the specific properties of $G$ crucially depend on the parameter choice, there are some general features that can be interpreted in the present context. Most importantly, at the bouncing time $t=t_{b}$, $G(t)$ actually diverges towards positive infinity from the right and to negative infinity from the left. The latter violates our requirement that $G(t)>0$, our assumption about smoothness as well as the condition $B(k)>-\frac{3}{4}\frac{1}{G(k)^{3}}\left(\frac{\partial G(k)}{\partial k}\right)^{2}$. Due to the dependence of $\phi$, $A^{2}$ and $V$ on $G$, this affects the whole STT. The scalar potential cannot be bounded from below, and, since $G(t)$ falls off to zero for very early and late times, $\phi$ diverges at infinity. Even worse, $\phi$ is given as the logarithm of $G$, which is not a real number for negative arguments. Overall, while the ELFP indeed allows us to arrive at a STT in which $\phi$ satisfies its EOM, the criteria of the equivalence are violated and it is not possible to define a consistent STT corresponding to this particular SD model.
@@ IV.3.3 Polytropic black hole with NEC
In the last example, we will deal with the polytropic black hole solution obtained in Ref. [136]. Due to the fact that this is a static, spherically symmetric solution, the only relevant variable is the radius $r$. Applying the NEC (79) to the effective energy-momentum tensor of SD gravity gives